An angle is 5 degrees less than twice another angle. Find the measures if they are supplementary angles.
Answer by rapaljer(4671)

Let x = one of the angles
2x-5 = second angle

Supplementary means that the angles add to 180 degrees.

x + 2x-5=180
3x - 5=180
3x=185
x=185/3=61.666...
2x-5=370/3-5=123.333...-5=118.333...

Check: 61.666... + 118.333... = 180 It checks!!

If anyone needs to contact me, especially about the website, my Email address is rapaljer@seminolestate.edu. I'll be glad to help you find an explanation on my website to help you with your algebra topic!

Dr. Robert J. Rapalje, Retired
Seminole State College of Florida
Altamonte Springs Campus